OpenStudy (anonymous):

using the expanding binomial theorem what is (3v-2u)^4

OpenStudy (phi):

do you know the "expansion" formula (from your notes) ?

OpenStudy (anonymous):

yes but not for using 2 variables

OpenStudy (phi):

the general form is nChooseK (a)^(n-k) (b)^k where k goes from 0 to n to use it, match up a with 3v and b with (-2u)

OpenStudy (anonymous):

how do i simplify it

OpenStudy (phi):

n is 4 so the first term (with k=0) is 4C0 * (3v)^(4-0) * (-2u)^0 4C0 is 1 (do you know how to do this part ? ) (3v)^4 means (3v) * 3v *3v*3v or 3^4 v^4 using a calculator or multiplying out , 3*3*3*3= 9*9= 81 we get 1 * 81*v^4 * (-2u)^0 anything to the zero power is 1, so the first term is 81 v^4

OpenStudy (phi):

do you follow how to do the first term ?

OpenStudy (anonymous):

ok so i just keep doing that to the rest of the terms right

OpenStudy (phi):

Let's do k=1: 4C1 * (3v)^(4-1) * (-2u)^1 can you simplify each part ?

OpenStudy (phi):

what is 4C1 ?

OpenStudy (anonymous):

4C1= (3v)^3 right

OpenStudy (phi):

no 4C1 or 4 choose 1 or \[\left(\begin{matrix}4 \\ 1\end{matrix}\right)= \frac{ 4! }{ 1! (4-1)! }\]

OpenStudy (phi):

or use your calculator (look for the key labeled nCr, for example)

OpenStudy (phi):

what is 4C1 ?

OpenStudy (anonymous):

i get 216v^2u^2 when i simplify the term

OpenStudy (phi):

what did you get for 4C1 ?

OpenStudy (anonymous):

4[(3v)^3 (-2)^1]

OpenStudy (phi):

4* 3^3 *(-1)^1 is what ?

OpenStudy (anonymous):

108

OpenStudy (phi):

I can't type today. 4* 3^3 *(-2)^1 is what ?

OpenStudy (anonymous):

-108 is what i got

OpenStudy (phi):

4*3*3*3*-2

OpenStudy (anonymous):

-216

OpenStudy (phi):

now we have 4C1 * (3v)^(4-1) * (-2u)^1 (4*3^3*-2) * v^3 * u^1 or -216 * v^3 * u^1 we leave out the multiply signs (the *'s) -216v^3 u (we leave out the exponent if it's 1... or we could put it in : -216v^3 u^1

OpenStudy (phi):

can you write down the formula when k=2 ?

OpenStudy (anonymous):

6[(3v)^2 (-2u)^2] which goes to 6*3*3*-2*-2

OpenStudy (phi):

yes, and for the letters?

OpenStudy (anonymous):

which is positive 216v^2u^2

OpenStudy (phi):

yes so far we have \[ 81v^4 -216v^3 u +216 v^2 u^2 + ...\] now k= 3

OpenStudy (phi):

the "letter part" follows a pattern.. we can write them down without thinking the V's exponent goes down by 1, the U's exponent goes up by 1... until we get to V^0 (which is 1)

OpenStudy (anonymous):

4[(3v)^1 (-2u)^3] which is 4*3*-2*-2*-2= -96vu^3 like that?

OpenStudy (phi):

yes \[ 81v^4 -216v^3 u +216 v^2 u^2 -96 v u^3+\] one last term k=4

OpenStudy (anonymous):

1[(3v)^0 (-2u)^4] which is 1*3^0*-2*-2*-2*-2= positive 16u^4

OpenStudy (phi):

and the whole expansion is \[ 81v^4 -216v^3 u +216 v^2 u^2 -96 v u^3+ 16u^4 \]

OpenStudy (phi):

the letters are obviously correct: v goes down from exponent 4 to 0 (last term v^0 is 1) and u goes up from exponent 0 to 4 the signs alternate (because -2^ odd number is negative and -2 to even exponent is +) so if we did not make a mistake with the numbers (I don't think we did), it looks good.

OpenStudy (anonymous):

ok thank you.. you taught me better than my teacher did